OpenStudy (anonymous):

how do you simply the limit algebraically: lim as x->0 (x^2)/(1-cos(x))

OpenStudy (anonymous):

answer please!!!!

OpenStudy (blockcolder):

\[\lim_{x \to 0} \frac{x^2}{1-\cos x}\cdot \frac{1+\cos x}{1+\cos x}=\lim_{x \to 0} \frac{x^2}{1-\cos^2x}\cdot \lim_{x \to 0}(1+\cos x)=1\cdot 2=2\]

OpenStudy (anonymous):

im starting to understand how you solve these, but how do you know the method that works? sometimes i try one way and the answer turns out to 0 and other ways it turns out to a real number (like this one), if i get the answer 0 should i just try the question another way to see if i get a different answer?

OpenStudy (blockcolder):

You get that with experience. :)

OpenStudy (anonymous):

alright, thank you!

OpenStudy (anonymous):

my friend told me to use the "L-hospital rule", would that be a viable method of doing these questions? The rule is to take the derivative of these functions.

OpenStudy (blockcolder):

Yes, but you have to make sure that you get 0/0 before doing it, and that there is no other way of doing it.

OpenStudy (anonymous):

0/0 as in when i just simply plug the lim x-> 0 in?

OpenStudy (blockcolder):

Yeah, like in this one.

OpenStudy (anonymous):

alright, thank you again!

OpenStudy (anonymous):

Apply L'Hospital's rule. Taking the derivative of the numerator over the derivative of the denominator twice \[ \frac {2x} {\sin (x)}\\ \frac 2 {\cos(x)} \] The limit of the second fraction is 2/1=2